The 1st, 5th and 10th terms of a linear sequence of a Geometric Progression. If the sum of the 2nd and 8th term is 30,find the first term
1
Expert's answer
2021-08-29T16:45:02-0400

In GP,

"1st = a"


"2nd =ar^2"

"5 th =ar^4"

"8 th =ar^7"

10th term "=ar^9"

Common Ratio (r) "=ar^4\/a= ar^9\/ar^4"

Common Ratio (r) "=r^4= r^5"

Common Ratio (r) = 1

2nd and 8th term = 30 ="ar+ ar^7=30"

"a(1)+ a(1)^7=30"

"a+a=30"

"2a = 30"

"a = 15"

Therefore, the first term is 15